OpenStudy (piercetheveil47):

The distance of planet Mercury from the Sun is approximately 5.8 times 10^7 kilometers, and the distance of Earth from the Sun is 1.5 times 10^8 kilometers. About how many more kilometers is the distance of Earth from the Sun than the distance of Mercury from the Sun? 4.3 Times 10^7 kilometers 9.2 Times 10^7 kilometers 9.2 Times 10^8 kilometers 5.7 Times 10^9 kilometers

OpenStudy (piercetheveil47):

@ganeshie8

OpenStudy (piercetheveil47):

is the answer A?

ganeshie8 (ganeshie8):

subtract both

ganeshie8 (ganeshie8):

\[1.5 \times 10^8 - 5.8\times 10^7\]

ganeshie8 (ganeshie8):

\[15 \times 10^7- 5.8\times 10^7\]

OpenStudy (piercetheveil47):

92000?

ganeshie8 (ganeshie8):

\[(15 - 5.8)\times 10^7\]

OpenStudy (piercetheveil47):

*92000000

ganeshie8 (ganeshie8):

its a bit hard to write/count so many zeroes

ganeshie8 (ganeshie8):

instead just put it in exponent form so that the reader knows how many 0's you mean lol

OpenStudy (piercetheveil47):

can you give me a min so i can chaco the answers and see which one is correct and then you can see if I'm correct please??

ganeshie8 (ganeshie8):

ok sure but we're almost doen

ganeshie8 (ganeshie8):

\[(15 - 5.8)\times 10^7 \] simplifies to \[9.2\times 10^7 \] end of story!

OpenStudy (piercetheveil47):

is it C??

OpenStudy (piercetheveil47):

ohhh!

OpenStudy (piercetheveil47):

okay!

OpenStudy (piercetheveil47):

thank you hahah

ganeshie8 (ganeshie8):

np :) remember to factor things and follow rules while simplifying these